OpenStudy (anonymous):

Find the circumference of a circle whose area is 452.16 square meters. (Use 3.1416 as the value of π.)

OpenStudy (anonymous):

75.83

OpenStudy (triciaal):

Area of circle, A = pi* r^2 = 452.16 Circumference of circle = 2* pi* r pi = 3.1416 using the formula for the area find r substitute in the formula for circumference